# Dark matter from phase transition generated PBH evaporation with gravitational waves signatures

### v1

@@ VI The combined spectra and detection prospects
If PBH are formed during the phase transition through Fermi-ball collapse, they can leave their indirect imprints in the GW spectrum. Considering the dominant sound wave contribution and using Eq. (17) and (18), we can relate the FOPT GW spectrum to the PBH mass and initial abundance as
$f_{\text{peak}}^{\rm PT}=\frac{3.65\times 10^{9}}{S^{1/3}}\beta_{\text{P}BH}^{1/3}\left(\frac{M_{\rm in}}{1~{}g}\right)^{-1/2}\text{Hz},$ (34)
$\Omega_{\rm peak}^{\rm PT}h^{2}=\frac{2\times 10^{-7}}{S^{4/3}}\beta_{\text{PBH}}^{2/3}\alpha^{-1/4}_{*}\eta_{\chi}^{-1}\left(\frac{\kappa_{\rm sw}\alpha_{*}}{1+\alpha_{*}}\right)^{2}.$ (35)
@@ VII Implications for DM phenomenology
BP5 $10^{4}$ 1.3 1.3 2.2 $2.7\times 10^{3}$ $1.6\times 10^{3}$ $1.4\times 10^{3}$ $1.4\times 10^{3}$ 206 0.13 0.79 $10^{-11}$
BP6 $10^{5}$ 1.4 1.3 2 $3.2\times 10^{4}$ $2.1\times 10^{4}$ $1.8\times 10^{4}$ $1.7\times 10^{4}$ 140 0.09 0.76 $10^{-10}$
@@
The dark sector fermion $\chi$ in our setup can not lead to the observed DM abundance, because of being trapped in the false vacuum. If we consider the dark sector to be purely gravitational in agreement with all experimental evidences, it can still be produced from PBH evaporation since the process of Hawking radiation is based solely on gravitational interactions. For the case of sub-dominant PBH ($\beta<\beta_{\rm crit}$), the DM relic density produced from PBH evaporation is found to be [^2] Here, for our demonstration, we consider the dark sector to consist of scalar particles. Similar results can be obtained for other particles depending on their degrees of freedom. Masina:2020xhk ; Bernal:2020kse ; Bernal:2020bjf ; JyotiDas:2021shi
@@
VIII Conclusion
@@ Acknowledgements
The work of DB is supported by the Science and Engineering Research Board (SERB), Government of India grant MTR/2022/000575. This work was supported by IBS under the project code, IBS-R018-D1. SJD would like to thank the support and hospitality of the Tata Institute of Fundamental Research (TIFR), Mumbai, where a portion of this project was carried out.

### v2

@@ VI The combined spectra and detection prospects
If PBH are formed during the phase transition through Fermi-ball collapse, they can leave their indirect imprints in the GW spectrum. Considering the dominant sound wave contribution and using Eq. (17) and (18), we can analytically relate the FOPT GW spectrum to the PBH mass and initial abundance approximately as
$f_{\text{peak}}^{\rm PT}\simeq\frac{5.33\times 10^{9}}{S^{1/3}}\beta_{\text{P}BH}^{1/3}\left(\frac{M_{\rm in}}{1~{}g}\right)^{-1/2}\text{Hz},$ (36)
$\Omega_{\rm peak}^{\rm PT}h^{2}\simeq\frac{1.42\times 10^{-7}}{S^{4/3}}\beta_{\text{PBH}}^{2/3}\alpha^{-1/4}_{*}\eta_{\chi}^{-1}\left(\frac{\kappa_{\rm sw}\alpha_{*}}{1+\alpha_{*}}\right)^{2}.$ (37)
@@ VII Implications for DM phenomenology
BP5 $10^{4}$ 1.3 1.3 2.2 $2.7\times 10^{3}$ $1.6\times 10^{3}$ $1.4\times 10^{3}$ $1.4\times 10^{3}$ 206 0.13 0.79 $10^{-12}$
BP6 $10^{5}$ 1.4 1.3 2 $3.2\times 10^{4}$ $2.1\times 10^{4}$ $1.8\times 10^{4}$ $1.7\times 10^{4}$ 140 0.09 0.76 $10^{-11}$
@@
The dark sector fermion $\chi$ in our setup can not lead to the observed DM abundance, because of being trapped in the false vacuum. However $\chi$ can still be produced from PBH evaporation Gehrman:2023qjn ; Kim:2023ixo or from the SM bath. Also, some small fraction of $\chi$ can cross from the false vacuum to the true vacuum Baker:2019ndr . While a detailed calculation of such trapping and crossing is beyond the scope of our present work, simple conservative estimates, shown in Appendix A justifies maximum trapping. In all such cases where $\chi$ appears in the true vacuum, the interaction of $\chi$ with the SM bath can play crucial role in deciding its final relic. Since we remain agnostic about the UV completion of the dark sector, we do not consider the possibility of $\chi$ being DM. The amount of $\chi$ produced from the above possible ways can finally decay into the SM bath, before the BBN epoch. As long as $\chi$ is long-lived enough and decay at a temperature $T_{\rm decay}\ll T_{*}$, the formation temperature of PBH, the phenomenology discussed here remain unaffected. Such a decay of $\chi$ can be introduced via explicit $U(1)_{D}$ global symmetry breaking higher dimensional operators like ¯$\overline{\ell}_{L}\tilde{H}\chi\Phi/\Lambda$ where $\ell_{L}$ denotes the lepton doublet of the SM. Since $\chi$ is lighter in the false vacuum, depending upon the UV completion of the dark sector, it is also possible to ensure DM decay in the true vacuum while preventing the same in the false vacuum due to kinematical restrictions.
If we consider the dark sector to be purely gravitational in agreement with all experimental evidences, it can still be produced from PBH evaporation since the process of Hawking radiation is based solely on gravitational interactions [^3] PBHs can also leave Planck remnants after evaporation, which can act as DM and have their imprints on GW. See Ref. Domenech:2023mqk .. For the case of sub-dominant PBH ($\beta<\beta_{\rm crit}$), the DM relic density produced from PBH evaporation is found to be [^4] Here, for our demonstration, we consider the dark sector to consist of scalar particles. Similar results can be obtained for other particles depending on their degrees of freedom. Masina:2020xhk ; Bernal:2020kse ; Bernal:2020bjf ; JyotiDas:2021shi
@@
VIII PBH from other mechanisms related to FOPT
Before concluding, we briefly comment on other FOPT origins of PBH and corresponding GW signatures. In other mechanisms without not relying on any dark sector asymmetry or Fermi ball collapse mechanism, PBH are formed when horizon-sized perturbations become more than the critical overdensity $\delta>\delta_{c}\sim 0.45$ leading to gravitational collapse Musco:2004ak . Such type of gravitational collapse as a consequence of FOPT have been studied in several works including Liu:2021svg ; Kawana:2022olo ; Hashino:2022tcs ; Gouttenoire:2023naa and references therein. During FOPT, nucleation may get delayed in some Hubble-sized region, and the late decay of the false vacuum increases the total energy density. In an expanding Universe, the radiation energy density keeps on decreasing, whereas in some regions the vacuum energy density remains constant. This leads to total energy density higher than the threshold overdensity and the mass inside such Hubble-sized region collapses to form PBH. After solving the Friedmann equation and radiation energy density evolution equations Liu:2021svg in late false vacuum decay region and background region, we can estimate the time, $t_{\rm PBH}$ when PBH formed and the initial PBH mass can be written as $M_{\rm PBH}=\frac{4\pi}{3}{\bf H}^{-3}(t_{\rm PBH})\rho_{c}$, where $\rho_{c}$ is critical energy density [^5] Similar relation is obtained for PBH formed through bubble collisionsJung:2021mku .. In such mechanisms, ultra-light PBH of mass $\sim 8\times 10^{8}$ g can be formed from FOPT with nucleation temperature of order $10^{13}$ GeV. Thus, if PBH is subdominant, the GW peak from FOPT will be in the ultra-high frequency regime $\sim 10^{7}$ Hz which is far outside the reach of near-future GW experiments [^6] There have been several proposed ideas to look for these ultra-high frequency gravitational waves. See Ref. Aggarwal:2020olq for a review.. For PBH domination, the GW peak from FOPT can get redshifted to a lower frequency (depending on $\beta_{\rm PBH}$) because of entropy dilution during PBH evaporation. For instance, considering the PBH mass mentioned above and taking $\beta_{\rm PBH}\sim\beta_{\rm max}$, we find the FOPT peak in the frequency range of order $10^{4}$ Hz, with a very small amplitude ($\Omega_{\rm peak}^{\rm PT}h^{2}\lesssim 10^{-25}$) because of the entropy dilution. Moreover, note that a detailed analysis requires calculating the energy fraction PBH $\beta_{PBH}$ through the probability of late false vacuum decay Kawana:2022olo . Therefore, the PBH formation mechanism adopted in this work offers more detection prospects irrespective of whether PBH dominates or remains subdominant.
IX Conclusion
@@ Acknowledgements
The work of DB is supported by the Science and Engineering Research Board (SERB), Government of India grant MTR/2022/000575 and CRG/2022/000603. The work of SJD was supported by IBS under the project code, IBS-R018-D1. SJD would like to thank the support and hospitality of the Tata Institute of Fundamental Research (TIFR), Mumbai, where a portion of this project was carried out.
Appendix A Particle trapping in false vacuum
Initially, all particles are in the false vacuum state. As the first order phase transition proceeds, particles gain mass in the true vacuum. The bubble of true vacuum expands and tries to capture the particles in the plasma. Due to the dynamical change of mass in the true vacuum, only sufficiently energetic particles are able to enter the bubble while conserving energy. Here, we show a simple quantitative analysis of particles getting trapped or reflected from bubble wall to false vacuum following the works of Chway:2019kft ; Gehrman:2023qjn . We can estimate the particle flux going to true vacuum in the bubble wall rest frame. The number of particles passing through the bubble wall along -z direction (considered for simplicity) per area $\Delta S$ in $\Delta t$ time can be written as
[…]
where, $r_{0}$ is bubble radius, $\Theta(-p_{z})$ is ensuring particles going inside and $\mathcal{T}(p)=\Theta(-p_{z}-m_{i})$ satisfied energy conservation inside bubble. Also, $g_{i}$ is degrees of freedom of particle with mass $m_{i}$ in true vacuum. Here, $f(p,x)$ is the distribution function of particles in false vacuum and it will be Bose-Einstein (BE) or Fermi-Dirac (FD) distribution. The distribution can be approximated to be Maxwell-Boltzmann distribution as $f(p,x)=e^{-\tilde{\gamma}(p+\tilde{v}p_{z})/T}$, where $\tilde{v}$ is the velocity of fluid or plasma with respect to bubble wall and $\tilde{\gamma}$ is corresponding Lorentz factor. The flux of particles in bubble wall frame can be calculated from above Eq. (41) as
$J_{w}=g_{i}T^{3}\left(\frac{\tilde{\gamma}(1-\tilde{v})m_{i}/T+1}{4\pi^{2}\tilde{\gamma}^{3}(1-\tilde{v})^{2}}\right)e^{-\frac{\tilde{\gamma}(1-\tilde{v})m_{i}}{T}}.$ (42)
Since, the bubble is expanding with a wall velocity $v_{w}$, the average number density inside true vacuum in global plasma frame can be expressed as
$n_{\rm in}=\frac{J_{w}}{\gamma_{w}v_{w}}.$ (43)
So, the fraction of particle which gets trapped in false vacuum can be written as
$F^{\rm trap}=1-\frac{n_{\rm in}}{n_{\rm eq}}$ (44)
where, $n_{\rm eq}$ is equilibrium number density of particles. In general, the velocity of fluid or plasma $\tilde{v}$ depends on fluid dynamics. From Eq. (44), it can be seen that $F^{\rm trap}$ is maximum ($\sim 1$) for fluid velocity $\tilde{v}$ O(0.1).